Advanced Scenario 6:

Matthew and Mary Donnelly Directions Using the tax software, complete the tax return, including Form 1040 and all appropri- ate forms, schedules, or worksheets . Answer the questions following the scenario .

Note: When entering Social Security numbers (SSNs) or Employer Identification Numbers (EINs), replace the Xs as directed, or with any four digits of your choice.

Interview Notes

• Matthew and Mary are married and want to file a joint return. They provided all the cost of keeping up the home and all of the support for their son Ryan .

• Ryan has no income and no filing requirement.

• Matthew retired and began receiving retirement income on April 1, 2018 . No distribu- tions were received prior to his retirement . Matthew selected a joint survivor annuity for these payments .

• Matthew was covered by Medicare all year . Mary and Ryan had minimum essen- tial healthcare coverage (MEC) through Mary’s employer until October 14th when she was laid off . Mary and Ryan did not have MEC all of November and December 2018 .

• Matthew and Mary stated if they are entitled to a refund, they want half of it depos- ited into their checking account and the other half deposited into their savings account . The checking account number is 123456789 and the savings account number is 987654321 .

18 . Ryan qualifies the Donnellys for the credit for other dependents. a . True b . False 19 .

19. The Donnellys must pay an individual shared responsibility payment because Mary and Ryan did NOT have healthcare coverage for each month of 2018 . a . True b . False

20. The net capital gain or loss reported on Schedule D, Capital Gains and Losses, is a gain of $638 . a . True b . False

21 . The combined age used to calculate the taxable portion of the pension using the Simplified Method is 129. a . True b . False

22 . None of Matthew’s Social Security income is taxable . a . True b . False

23 . The Donnellys can split their refund using Form 8888, Allocation of Refund (Including Savings Bond Purchases) . a . True b . False

24 . The total withholding on the tax return is $4,146 . a . True b . False
Answer #1

18. True.

19. True. The married couple who can claim a child or another individual as a dependent for federal income tax purposes is responsible for making the payment if the dependent does not have the coverage or an exemption.

20. True. 1088 - 450 = 638

21. False.

22. False. If you are married filing jointly and you lived with your spouse for any part of the tax year all your Social Security income is taxable.

23. True. The Donnellys can split their refund using Form 8888.

24. False.

Total withholding on the tax return is = 620 + 911.40 + 213.15 + 1935 + 1591 = 5270.55